killer:

Solve the inequality. Graph the solution. 3p+2≥−10 The solution is

jhonyy9:

like a first step subtract from both sides 2 so what will get ?

jhonyy9:

@supie any idea ?

supie:

So when i subtract 2 from both sides i got \(3p≥−12\) because \(3p+2−2≥−10−2\) -10-2=12 so \(3p≥−12\) then divide both sides by 3 ???

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie So when i subtract 2 from both sides i got \(3p≥−12\) because \(3p+2−2≥−10−2\) -10-2=12 so \(3p≥−12\) then divide both sides by 3 ??? \(\color{#0cbb34}{\text{End of Quote}}\) good idea

jhonyy9:

continue pls.

jhonyy9:

@killer ok ?

supie:

Ok so \(\LARGE\frac{3p}{3}≥\frac{-12}{3}\) \(\cancel{\frac{3p}{p}}≥\frac{-12}{3}\) (3p/p cancels) so all that is lef tis \(\frac{-12}{3}\) so that needs to be simplified \(\frac{-12}{3}=?\) \(p≥?\)

supie:

left is * ._.

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Ok so \(\LARGE\frac{3p}{3}≥\frac{-12}{3}\) \(\cancel{\frac{3p}{p}}≥\frac{-12}{3}\) (3p/p cancels) so all that is lef tis \(\frac{-12}{3}\) so that needs to be simplified \(\frac{-12}{3}=?\) \(p≥?\) \(\color{#0cbb34}{\text{End of Quote}}\) attention pls on the left side 3 cancel by 3 not p by p

jhonyy9:

3p/3 not 3p/p

supie:

Oh that's what I meant.

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Ok so \(\LARGE\frac{3p}{3}≥\frac{-12}{3}\) \(\cancel{\frac{3p}{p}}≥\frac{-12}{3}\) (3p/p cancels) so all that is lef tis \(\frac{-12}{3}\) so that needs to be simplified \(\frac{-12}{3}=?\) \(p≥?\) \(\color{#0cbb34}{\text{End of Quote}}\) attention pls on the left side 3 cancel by 3 not p by p \(\color{#0cbb34}{\text{End of Quote}}\) on the left side first you ve wrote 3p/3 but you continue with 3p/p

jhonyy9:

@supie

supie:

Yes i know. I meant \(\frac{3p}{3}\) but I accidently did \(\frac{3p}{p}\)

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Ok so \(\LARGE\frac{3p}{3}≥\frac{-12}{3}\) \(\cancel{\frac{3p}{p}}≥\frac{-12}{3}\) (3p/p cancels) so all that is lef tis \(\frac{-12}{3}\) so that needs to be simplified \(\frac{-12}{3}=?\) \(p≥?\) \(\color{#0cbb34}{\text{End of Quote}}\) attention pls on the left side 3 cancel by 3 not p by p \(\color{#0cbb34}{\text{End of Quote}}\) on the left side first you ve wrote 3p/3 but you continue with 3p/p \(\color{#0cbb34}{\text{End of Quote}}\) @killer did you get the answer now ?

jhonyy9:

questions ?

jhonyy9:

@supie how you graph the solution ?